Can you please help me with this linear equation problem?
The difference between two numbers is 34. The larger exceeds 3 times the smaller by 4. Find the numbers.
Thanks


Answer by est_qed(2)

Solution:
y - x = 34.
Suppose y > x, then
y = 3x + 4.
Hence
3x + 4 = 34 + x
or
2x = 30.
That is
x = 15 and y = 34 + 15 = 49.